Find two square roots for the complex number. Write your answer in standard form. -32-321/3
Answer #1

ez (say) - 32 - 32iB (ariy) = x - yr + 2 ayi comparing real of imaginary parts, x-yu = -32 2 2xy = -3253 that know (2ľ = 12²If you have any doubt, please comment and I will try my best to explore this to you. Thank you.